FLUORESCENCE-DIP INFRARED SPECTROSCOPY OF METHOXY RADICAL

\underline{JINJUN LIU}, LILY ZU, WEIDONG ZHOU, VLADIMIR LOZOVSKY, PATRICK DUPRÉ, C. BRADLEY MOORE AND TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

The fluorescence-dip infrared (FDIR) spectrum of methoxy radical (CH_{3}O) has been recorded in the region 2875-2980 cm^{−1}. Methoxy radical was produced by the photolysis of methyl nitrite (CH_{3}ONO) in a pulsed free-jet expansion. The laser induced fluorescence (LIF) spectrum obtained was characterized by a rotational temperature of ∼2K. In the FDIR experiment, the UV excitation laser was tuned to the rotational transition, N′=1, K′=0 ← J′′=3/2, P′′=3/2 of the Ã 2A_{1} ← X̃ ^{2}E_{3/2}, 0-0 band. Two IR sources were used to deplete the ground-state level: difference frequency mixing and an optical parameter oscillator (OPO). The FDIR spectrum is detected as dips in fluorescence as a function of IR frequency. The correlation between the FDIR spectrum, the IR absorption spectrum,^{a} the stimulated emission pumping (SEP) spectrum and the dispersed fluorescence (DF) spectrum will be discussed.

aJia-xiang Han, Yu. G. Utkin, Hong-bing Chen, L. A. Burns, and R. F. Curl, J. Chem. Phys. \underline{117}, 6538, (2002)